OpenStudy (anonymous):

Determine whether the points P1, P2, and P3 lie on the same line. P1(6, 9, 7), P2(9, 2, 0), P3(0,−5,−3)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

please help me learn to do this. thanks

OpenStudy (anonymous):

@Loser66

OpenStudy (baru):

do you know how to write the equation of a line given two points?

OpenStudy (anonymous):

yes but this is in 3D not 2D

OpenStudy (baru):

yes, you will have to write a parametric equation

OpenStudy (anonymous):

x=x0+at , y=y0+bt , z=z0+ct the form for the equation of the line in 3D

OpenStudy (anonymous):

i need to find the vector parallel to the line? i think but how shd i go about doing so

OpenStudy (baru):

just take the first two points P1 and P2 (P2-P1) is the vector connecting P2 and P1

OpenStudy (anonymous):

yes i did that then do i take point (p3-p2)? p2-p1= (3,-7,-7)

OpenStudy (baru):

now with (3,-7,-7) (this is a vector paralell to the line) write the equation of the line. Then check is the point P3 satisfies the equation

OpenStudy (anonymous):

the equation of the line would be x=6+3t , y=9-7t , z=7-7t

OpenStudy (anonymous):

i dont think i know how to check to see if p3 satisfy the equation ...

OpenStudy (baru):

just match the x co-ordinate and find a value for t P3 has x=0 that means t=-2 but when t=-2 y=23 z=21 so P3 is not a point on the line

OpenStudy (anonymous):

o ok thanks i understand now

OpenStudy (baru):

come to think of it... there are easier methods to do this

OpenStudy (anonymous):

there is? please share it ...hanks

OpenStudy (baru):

have you learnt the "cross product"?

OpenStudy (anonymous):

yes

OpenStudy (baru):

take the vectors (P2-P1) and(P3-P1) take their cross product, if they are parallel, it will be zero

OpenStudy (baru):

if those two vectors are parallel, then the three points lie on the same line

OpenStudy (anonymous):

p2-p1=(3,-7,-7) p3-p1=(-6,-14,-10) cross product . i(70-98)-j(-30-42)+k(-42-42) =-28i+72j-84k

OpenStudy (baru):

yep, so same conclusion, they dont lie on the same line

OpenStudy (anonymous):

ok thanks

OpenStudy (baru):

:)